Oxalic acid dihydrate is a solid, diprotic acid that can be used in the laboratory as a primary standard. Its formula is H2C2O4•2H2O.

A student dissolves 0.465 grams of H2C2O4•2H2O in water and titrates the resulting solution with a solution of potassium hydroxide of unknown concentration. If 21.0 mL of the potassium hydroxide solution are required to neutralize the acid, what is the molarity of the potassium hydroxide solution?

___________ M
1
Expert's answer
2019-03-21T06:24:53-0400
"H_2C_2O_4\cdot 2H_2O + 2KOH = K_2C_2O_4 + 4H_2O"

"M(H_2C_2O_4\cdot H_2O) = 126.08 \frac{g}{mol}"


"n=\frac{m}{M}"


"n(C_2H_2O_4\cdot 2H_2O) = \frac{0.465}{126.08} = 0.00369 mol"


According to equation mole ratio "n(C_2H_2O_4\cdot 2H_2O):n(KOH) = 1:2" , then


"n(KOH) = 2\times 0.00369 mol = 0.00738 mol"


"c = \frac{n}{V}"


"c = \frac{0.00738 mol}{0.0210L} = 0.351 \frac{mol}{L} = 0.351 M"


Answer: 0.351 M